Inventory Analysis

The following data were extracted from the income statement of Keever Inc.:

Current Year Previous Year
Sales $1,430,800 $1,498,000
Beginning inventories 108,884 100,868
Cost of goods sold 715,400 832,200
Ending inventories 98,484 108,884

a. Determine for each year (1) the inventory turnover and (2) the number of days' sales in inventory. Round interim calculations to the nearest dollar and the final answers to one decimal place. Assume 365 days a year.

Current Year Previous Year
1. Inventory turnover
2. Number of days' sales in inventory days days
Answer #1
FORMULA CURRENT YEAR PREVIOUS YEAR
AVERAGE INVENTORY =BEGINNING INVENTORY + ENDING INVENTORY / 2 = 108884 + 98484 / 2 = 100868 + 108884 / 2
103684 104876
DAY'S SALES IN INVENTORY = AVERAGE INVENTORY / COGS x 365 = 103684 / 715400 x 365 = 104876 / 832200 x 365
52.9 DAYS 45.9 DAYS
INVENTORY TURNOVER =SALES / AVERAGE INVENTORY =1430800 / 103684 =1498000 / 104876
13.8 TIMES 14.3TIMES
ALTERNATIVE = COGS / AVERAGE INVENTORY =715400 / 103684 =832200 / 104876
6.9 TIMES 7.9 TIMES